I am having issues with these two questions
IN a recent basketball game the boston celtics scored 99 points They scored three times as many field goals as free throws.They also made 12 three point buckets. How many field goals did they make, how many free throws did they make. (Field goal is worth 2 points) (Free throw is worth 1)
A running back in Wheaton college football game catches a pass and runs for the goal line 80 yards away. He ran a 50 yard dash In six seconds. If he runs at 80% of his speed in the 50 yard dash, how long will it take him to reach the goal line

Answer by rothauserc(4718)

1)let x be the number of free throws
3(2x) +x + 12(3) = 99
6x +x +36 = 99
7x = 63
x=9
free throw points = 54 and number of free throws = 54/2 = 27
2) rate times time = distance
rate * 6 = 50
rate = 50/6 = 8.33 yards/sec
now .80 * 8.33 = 6.66 yards/sec for new rate
6.66 * time = 80
time to reach goal line = 80/6.66 = 12.01 seconds